OpenStudy (anonymous):

A denim manufacturer has determined that for every 13 pairs of jeans produced, there are 3 defects in workmanship. If the manufacturer produces 14,300 pairs of jeans this week, what is the best prediction for how many defects there will be?

OpenStudy (anonymous):

3300

OpenStudy (anonymous):

3/13=x/14,300 x=(14,300x3)/13 x=3,300